two positive numbers have a diference of d.Their reciprocals have a difference of r.Find an expression for their product,xy,in terms of d and r.
Answer by stanbon(75887)

two positive numbers have a diference of d.Their reciprocals have a difference of r.Find an expression for their product,ab,in terms of d and r.
----------------------
Equations:
a - b = d
1/a - 1/b = r
--------
Modify the 2nd equation to get ab:
(b-a)/ab = r
---
So, ab = (b-a)/r
---
But b-a = -d
----
Therefore: ab = -d/r
=========================
Cheers,
Stan H.
=======================